In the electrolysis of melted MgCl2, an electric current of 52115 Coulomb is used, the mass of Mg produced is ...



1
Expert's answer
2021-02-22T06:09:57-0500

Solution:

Electrolysis of the molten MgCl2:

Mg2+ + 2e- → Mg (at the (-)cathode).

2Cl- - 2e- → Cl2 (at the (+)anode).

The overall reaction is:

MgCl2(l) → Mg(s) +   Cl2(g)


By Faraday's first law

w = z × Q

where:

z = Electrochemical equivalent = Atomic mass / (Electrons × 96500)

Q = Coulombs of electricity


Hence,

Mass of Mg = (M × Q) / (ne × F)

Mass of Mg = (24 × 52115) / (2 × 96500) = 6.48 g

Mass of Mg = 6.48 g

Answer: 6.48 grams of Mg is produced.